Here’s a very short sample BibTeX document. Among the interesting books and papers I have read in the course of this work is Halzen and Martin’s Quarks and Leptons [1], and the paper Can one probe the structure of the pomeron? by Ellis and Ross [2].

And here the sample document ends.

Bibliography

[1] F. Halzen and A. D. Martin, QUARKS AND LEPTONS: AN INTRO-DUCTORY COURSE IN MODERN PARTICLE PHYSICS (Wiley, New York, USA, 1984).

[2] J. Ellis and G. G. Ross, Phys. Lett. B384, 293 (1996), hep-ph/9604360.